Circular No. 7547

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000fo IN PGC 70148 M. Schwartz, Cottage Grove, OR, reports the discovery by LOTOSS

(cf. IAUC 7514) of an apparent supernova (mag ∼ 16.5) on unfiltered images taken with the 0.5-m Tenagra III automatic telescope on Dec. 21.1 and 22.0 UT. The new object is located at α = 22^{h}58^{m}25^{s}.34, δ = +26^{o}38′26′′.6 (equinox 2000.0), which is 5′′.5 west and 5′′.8 north of the nucleus of PGC 70148. An image of the same field taken with the Katzman Automatic Imaging Telescope (cf. IAUC 6627, 7126) under poor conditions on Dec. 13.1 already showed a hint of SN 2000fo, while an image of the field taken with the 0.5-m Tenagra III automatic telescope on Dec. 6.1 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVAE 2000fm, 2000fn, AND 2000ew A. V. Filippenko, R. Chornock, and M. Modjaz, University of Califor-

nia at Berkeley, report that inspection of uncalibrated CCD spectra (range 330–1000 nm) of SN 2000fn (IAUC 7546) obtained on Dec. 21 UT with the Shane 3-m reflector at Lick Observatory reveals that the object is a type-Ib supernova ∼ 1 month past maximum brightness. Lower-quality spectra reveal that SN 2000fm (IAUC 7545) is probably a very young type-II supernova; relatively weak Hα emission is visible on a blue, featureless continuum. High-quality spectra of SN 2000ew (IAUC 7530) indicate that it is actually of type Ic (not type Ia, as stated on IAUC 7532).

RZ LEONIS J. A. Mattei, American Association of Variable Star Observers, reports

that this WZ Sge-type cataclysmic variable is having a bright outburst (possibly a superoutburst), as indicated by the following visual magnitude estimates (S = R. Stubbings, Drouin, Victoria, Australia; D = P. Dubovsky, Podbiel, Slovakia): Dec. 4.404 UT, [14.0 (S); 7.165, [14.0 (D); 9.729, [12.0 (S); 20.662, 12.1 (S); 20.700, 12.1 (S); 20.720, 12.2 (S). High-speed photoelectric or CCD photometry to search for superhumps is strongly recommended. The most recent outburst recorded in the AAVSO International Database was in 1990, when the star was in outburst for at least 5 days between Oct. 26 and 31, reaching visual mag 12.8.

2000 December 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green